# Magnetic field dynamics in presence of Hall conductivity, and thermodiffusion

### v1

@@ Abstract
This anisotropy of kinetic coefficients in presence of a magnetic field is represented by so-called Hall currents, which appear in a collisional medium due to action of the Lorentz force on the charged particles between collisions. In many papers the Hall currents had been considered in different approximate approaches. We derive equations, describing dynamics of the magnetic field in presence of Hall currents, using a standard electrodynamic consideration. We consider collisional media, and take into account a temperature gradients, which create thermodiffusional electric current, in presence of the Hall component. The influence of the Hall currents on the magnetic field structure and damping is considered in simple models. In presence of thermodiffusion the condition for creation of the seed magnetic field in the non-magnetized media is found, which is needed for the action of the mechanism, known as ”Biermann battery”.
@@ 1 Introduction
In presence of a magnetic field kinetic coefficients, describing charged particles, became anisotropic. It is related to the plasma, and to the metals. This anisotropy represents appearance of so-called Hall currents. They appear in a collisional medium due to action of the Lorentz force on the charged particles between collisions. There is an ample literature where different aspects of the Hall currents were investigated, in particular, their influence on the magnetic field structure and damping. The most striking feature of the Hall currents is their property of preserving entropy, and absence of a heat production. Therefore the magnetic field damping due to Hall currents may happen only indirectly, by changing the magnetic field structure. In many papers (see i.g. [1],[2], [3],[4],[5],[6]) the account if the Hall current had been done with some approximations, including works with consideration of GR effects [7].
Here we derive equations, using exact electrodynamic consideration, describing dynamics of the magnetic field in presence of Hall currents. We consider collisional media, like fully ionized plasma in the space and in the laboratory. We have taken into account a temperature gradients, which create thermodiffusional electric current, in presence of the Hall component.
In presence of thermodiffusion the condition for creation of the seed magnetic field in the non-magnetized media is found, which is needed for the action of the mechanism, known as ”Biermann battery” [8, 9]. The influence of the Hall currents on the magnetic field structure and damping is considered in simple models.
1.1 MHD equations in presence of Hall conductivity
@@
${\bf j}=\sigma_{ij}\{{\bf E}+\frac{1}{c}{\bf[v\times B]}\}.$ (3)
Anisotropic conductivity tensor $\sigma_{ij}$ has a relatively simple form in the first Chapman-Enskog approximation [11]. Finally the expression in (3) may be written in the form [3]-[6]:
@@
Here $\omega=(eB/m_{e}c)$ is the electron Larmor frequency, $\tau_{e}$ is a time between electron-ion collision, in Lorentz approximation; $n_{e}$, $n_{N}$ are concentrations of the electrons and nuclei with atomic number Z, $\Lambda$ is a Coulomb logarithm, ${\bf E}$, ${\bf E^{\prime}}$ are electric fields in the laboratory, and comoving coordinate system, respectively.
${\bf E^{\prime}}={\bf E}+\frac{1}{c}{\bf[v\times B]}.$ (6)
In Cartesian coordinates $(x,y,z)$ with
${\bf E^{\prime}}=(E^{\prime}_{x},\,E^{\prime}_{y},\,E^{\prime}_{z});\qquad{\bf B}=(B_{x},\,B_{y},\,B_{z}),$ (7)
the generalized Ohm law (4) is represented by the following three equations:
$\frac{1+\omega^{2}{\tau_{e}}^{2}}{\sigma_{0}}j_{x}=E^{\prime}_{x}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}(B_{x}E^{\prime}_{x}+B_{y}E^{\prime}_{y}+B_{z}E^{\prime}_{z})B_{x}-$
$\frac{\omega\tau_{e}}{B}(B_{y}E^{\prime}_{z}-B_{z}E^{\prime}_{y}),$ (8)
$\frac{1+\omega^{2}{\tau_{e}}^{2}}{\sigma_{0}}j_{y}=E^{\prime}_{y}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}(B_{x}E^{\prime}_{x}+B_{y}E^{\prime}_{y}+B_{z}E^{\prime}_{z})B_{y}-$
$\frac{\omega\tau_{e}}{B}(B_{z}E^{\prime}x-B_{x}E^{\prime}_{z}),$ (9)
$\frac{1+\omega^{2}{\tau_{e}}^{2}}{\sigma_{0}}j_{z}=E^{\prime}_{z}+\frac{\omega^{2}{\tau_{e}}^{2}}{B^{2}}(B_{x}E^{\prime}_{x}+B_{y}E^{\prime}_{y}+B_{z}E^{\prime}_{z})B_{z}-$
$\frac{\omega\tau_{e}}{B}(B_{x}E^{\prime}_{y}-B_{y}E^{\prime}_{x}).$ (10)
For deriving a dynamic equation for the magnetic field ${\bf B}$ evolution in conditions of anisotropic conductivity we need to solve the system of the above linear equations for finding a dependence ${\bf E^{\prime}=E^{\prime}(j,B)}$, for substituting it in the first equation (1). We have found the solution in the following form:
$E^{\prime}_{x}=\frac{1}{\sigma_{0}}[j_{x}-\frac{\omega\tau_{e}}{B}(B_{z}j_{y}-B_{y}j_{z})],\qquad\qquad$ (11)
$E^{\prime}_{y}=\frac{1}{\sigma_{0}}[j_{y}-\frac{\omega\tau_{e}}{B}(B_{x}j_{z}-B_{z}j_{x})],\qquad\qquad$ (12)
$E^{\prime}_{z}=\frac{1}{\sigma_{0}}[j_{z}-\frac{\omega\tau_{e}}{B}(B_{y}j_{x}-B_{x}j_{y})].\qquad\qquad$ (13)
With account of (2), this solution is written in the vector form as:
@@
This equation had been written in [10],[18], where a third-party E.D.S. was also included.
Substituting (14) into (1), with account of (6), we obtain the equation for the evolution of a magnetic field at anisotropic electric conductivity due to Hall current, in the form [2],[3],[19],[5],[6]:
${\frac{\partial{\bf B}}{\partial t}={\bf\nabla}\times[{\bf v\times B}]+\frac{c^{2}}{4\pi\sigma_{0}}{\bf\nabla^{2}B}+\frac{c^{2}\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf\nabla}\times[{\bf B\times[\nabla\times B]}]}.$ (15)
@@
The equation (15) is describing the evolution of the magnetic field in the medium moving with center mass velocity $\bf v$ relative to the laboratory frame. In presence of a magnetic field $\bf B$ it has an anisotropic conductivity, with the Hall component represented by the last term in (15).
It is interesting to compare (15) with the equation (39) from the paper [1], where the first term to the left has the same dependence on $\bf B$, as the last term in (15), but has a different coefficient. The authors of [1] have called it as ”Hall drift”. Actually this term is connected with Lorentz transformation of the electrical field E from the laboratory frame to the co-moving one [10]. The Hall effect is connected with the anisotropic electrical conductivity in the conducting medium in presence of the magnetic field, which is included in the generalized Ohms law [11], given here in Eq.(4). In [1] the equation (39) is applied to the metallic static body with the electric current $\bf j$, which is obtained from the Lorentz transformation equation by identifying ${\bf j}$= $e{\bf v}$.
This procedure is not correct, because the center mass ${\bf v}$, connected with the moving frame is actually zero, and the electric current should be found from the Ohms law, like in Eq.(4). The model in [1] is almost the same as in the case of the plasmic body, because the electrical current is produced mainly by the electron motion in the medium moving with the velocity ${\bf v}$, which is zero in this case.
The difference between plasmic and metallic body consist of difference in the scalar electrical conductivity $\sigma_{0}$, and in different ways of supplying a static state. For the metal bar the motion is described by the mechanic law, and a plasmic liquid body is governed by the MHD equations. According to (15), in the ideal plasma with infinite $\sigma_{0}$, there is no damping of the magnetic field, no influence of the Hall currents on its motion and in the co-moving frame the electric field is zero. In [1] the so-called ”Hall drift” term remains non-zero in the ideal plasma, what leads to artificial behavior of the magnetic field in it.
2 Equations in the presence of thermodiffusion, and external electro-moving force (EMF)
With account of thermodiffusion and external electrical field ${\bf E^{{}^{\prime\prime}}}$, the Ohm’s law has the following form, when using the same approximation as in the derivation of (4), and remaining only the term with a temperature gradient ${\bf\nabla}T$ in the diffusion vector ${\bf d}$, see [15]:
@@ 2 Equations in the presence of thermodiffusion, and external electro-moving force (EMF)
The linear equations, connecting the components of the electrical current $j_{i}$, electrical field in the co-moving frame $E^{\prime}_{i}$ and temperature gradient $\frac{\partial T}{\partial x_{i}}$ are written as ($i=x,y,z$):
@@
With account of (2), this solution is written in the vector form as:
@@
${\bf E}=-\frac{1}{c}{\bf[v\times B]}-{\bf E^{{}^{\prime\prime}}}-\frac{\lambda_{0}}{\sigma_{0}}{\bf\nabla}T+\frac{c}{4\pi\sigma_{0}}{\bf\nabla\times B}-\frac{c\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf B\times[\nabla\times B]}.$ (26)
In presence of thermodiffusion and external electrical field $\bf E^{{}^{\prime\prime}}$ we obtain, after applying the curl procedure to the equation (26), and using (1), the following equation:
$\frac{\partial{\bf B}}{\partial t}={\bf\nabla}\times[{\bf v\times B}]+c{\bf\nabla\times E^{{}^{\prime\prime}}}+c\nabla\biggl{(}\frac{\lambda_{0}}{\sigma_{0}}\biggr{)}\times{\nabla T}\qquad$ (27)
@@
The Eq.(27) is valid for variable values of the parameters $\lambda_{0}$ and $\sigma_{0}$. At constant value $\lambda_{0}/\sigma_{0}=3k/2e$ in the fully ionized Lorenz plasma, the Eq.(27) does not contain the temperature terms. It means that in this conditions thermodiffusion is not participating directly in the magnetic field evolution. In this case the Eq.(27) coincides with the similar equation from [10].
At variable transport coefficients the temperature term in the Eq.(27) may be non-zero even at initially zero values of ${\bf B}$ and ${\bf E}$, what leads to creation of a seed magnetic field in the non-magnetized media. At the function $\lambda_{0}/\sigma_{0}=f(\rho,T)$, the seed magnetic field is created at non-zero value of
@@
It is representing the equation for Biermann battery [8, 9], which is derived here by the solution of Boltzmann kinetic equation.

### v2

# Magnetic field dynamics in presence of thermo-diffusion and Hall conductivity

@@ Abstract
Anisotropy of kinetic coefficients in presence of a magnetic field is represented by Hall currents, which appear in a collisional medium due to action of the Lorentz force on the charged particles between collisions. We derive equations, describing dynamics of the magnetic field in presence of thermo-diffusion with Hall currents, using a standard electrodynamic consideration. The influence of the Hall currents, at presence of thermo-diffusion, on the magnetic field structure is considered in simple models. The equation is derived, which includes additional term for a seed magnetic field creation in the non-magnetized plasma, due to thermo-diffusion. This equation describes the action of the seed magnetic field creation by the mechanism known as ”Biermann battery”.
@@ 1 Introduction
In presence of a magnetic field, kinetic coefficients, describing charged particles, became anisotropic. It is related to plasma, as well as to metals. Such anisotropy represents appearance of Hall currents, which appear in a collisional plasma, due to action of the Lorentz force on the charged particles between collisions. The characteristic feature of Hall currents is their property of preserving entropy, and absence of a heat production. Therefore the influence of Hall currents on a magnetic field damping may happen only indirectly, by changing the magnetic field structure.
Transport coefficients in the non-degenerate magnetized plasma had been found in [11], [12], using solution of the Boltzmann kinetic equation by Chapmen-Enskog method [10]. Thermo-diffusion and Hall currents had been taken into account in these papers, together with analysis of different physical effects. The equation for the excitation of a seed magnetic field in the non-magnetized plasma, obtained there, is connected with a pressure gradient term in the diffusion vector $\bf d_{e}$} [10], [12]
${\bf d_{e}}=\frac{e}{kT}\large[{\bf E}+\frac{1}{c}[{\bf v}\times{\bf B}]\large]+\frac{{\bf\nabla}P_{e}}{P_{e}}=\frac{e}{kT}\large[{\bf E}+\frac{1}{c}[{\bf v}\times{\bf B}]+\frac{{\bf\nabla}P_{e}}{en_{e}}\large]$ (1)
We have derived here the equation for the excitation of a seed magnetic field in presence of thermo-diffusion, and have obtained the additional term in the equation for seed field excitation. The creation of a seed magnetic field is a very important astrophysical problem. The presence of magnetic field everywhere in the universe is explained by different mechanisms of its amplification, but these processes could happen only after creation of initial seed field in the initially non-magnetized universe,
Development of a laboratory astrophysics is aimed to model astrophysical events in a terrestrial laboratory. MHD simulations with Hall current had been performed in [37], to simulate the laboratory experiments. Numerical calculations of the magnetic field evolution in neutron star crust had been done in [4]-[5] in presence of the Hall effect. Here similar equations are derived in presence of a temperature gradient, external action, and thermo-diffusion.
We have separated magnetic and electric fields on internal and external one, depending on their origin: internal electrical currents and electrical charge distribution, and external source. Those could be the electrical field produced by battery, accumulators etc., and magnetic field originated by external currents, or external non-vortex magnetic field.
In presence of thermo-diffusion, the additional channel for creation of the seed magnetic field in the non-magnetized media is found, which is needed for the action of the mechanism, known as ”Biermann battery” [7, 8].
Simple models are considered for illustration of Hall current’s influence on magnetic filed structure, including the action of thermo-diffusion.
2 MHD equations in presence of Hall conductivity
@@
${\bf j}=\sigma_{ij}{{\bf E^{\prime}}}.$ (4)
Anisotropic conductivity tensor $\sigma_{ij}$ has a relatively simple form in the first Chapman-Enskog approximation, $\bf E^{\prime}$} is the the electrical field in the co-moving coordinate system. The expression in (4) may be written in the form [3]-[6]:
@@
Here $\omega=(eB/m_{e}c)$ is the electron Larmor frequency, $\tau_{e}$ is a time between electron-ion collision, in Lorentz approximation; $n_{e}$, $n_{N}$ are concentrations of the electrons and nuclei with atomic number Z, $\Lambda$ is a Coulomb logarithm, $\bf E$}, $\bf E^{\prime}$} are electric fields in the laboratory, and comoving coordinate system, with account of the pressure gradient term, respectively . The comoving system is moving with velocity $\bf v$} relative to the laboratory system, together with matter, so that [9],[12]
${\bf E^{\prime}}={\bf E}+\frac{1}{c}{\bf[v\times B]}+\frac{{\bf\nabla}P_{e}}{en_{e}}.$ (7)
From equations (2)-(5) follows the expression for $\bf E^{\prime}$}, as a function of $\bf j,\,B$}, with account of Hall currents, in the form [9],[17]
@@
Substituting (8) into (2), with account of (7), leads to the equation for the evolution of a magnetic field [12], [2],[3],[18],[5],[6], at anisotropic electric conductivity due to Hall current, as
$\frac{\partial{\bf B}}{\partial t}={\bf\nabla}\times[{\bf v\times B}]+\frac{c^{2}}{4\pi\sigma_{0}}{\bf\nabla^{2}B}$ (9)
$+\frac{c^{2}\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf\nabla}\times[{\bf B\times[\nabla\times B]}]-\frac{ck}{en_{e}}[{\bf\nabla}n_{e}\times{\bf\nabla}T].$
@@
The equation (9) is describing the evolution of the magnetic field in the medium moving with center mass velocity $\bf v$ relative to the laboratory frame.
It is interesting to compare (9) with the equation (39) from the paper [1], where the first term to the right has the same dependence on $\bf B$, as the second term from the end in (9), but has a different coefficient. The authors of [1] have called it as ”Hall drift”. Actually this term is connected with Lorentz transformation of the electrical field E from the laboratory frame to the co-moving one [9]. The Hall effect is connected with the anisotropic electrical conductivity in the conducting medium in presence of the magnetic field, which is included in the generalized Ohms law [10], given here in Eq.(5). In [1] the equation (39) is applied to the metallic static body with the electric current $\bf j$, which is obtained from the Lorentz transformation equation by identifying $\bf j$}= $e{\bf v}$. This procedure is not correct, because the center mass $\bf v$}, related to the moving frame, is actually zero. According to (9), in the ideal plasma with infinite $\sigma_{0}$, and absence of the pressure gradient term, there is no damping of the magnetic field, no influence of the Hall currents on its behavior, and in the co-moving frame the electric field is zero. In [1] the so-called ”Hall drift” term remains non-zero in the ideal plasma, what leads to artificial behavior of the magnetic field in it.
3 Equations in the presence of thermo-diffusion, and external electro-moving force (EMF)
With account of thermo-diffusion and external electrical field $\bf E^{{}^{\prime\prime}}$}, see [9], [10],[11],[12], [13], the Ohm’s law has the following form:
@@ 3 Equations in the presence of thermo-diffusion, and external electro-moving force (EMF)
The linear equations, connecting the components of the electrical current $j_{i}$, electrical field in the co-moving frame $E^{\prime}_{i}$, external electrical field $E^{\prime\prime}_{i}$, and temperature gradient $\frac{\partial T}{\partial x_{i}}$ are written as ($i=x,y,z$):
@@
With account of (3), (7), this solution is written in the vector form as:
@@
${\bf E}=-\frac{1}{c}{\bf[v\times B]}-{\bf E^{{}^{\prime\prime}}}-\frac{\lambda_{0}}{\sigma_{0}}{\bf\nabla}T+\frac{c}{4\pi\sigma_{0}}{\bf\nabla\times B}$ (20)
$-\frac{c\,\omega\tau_{e}}{4\pi\sigma_{0}B}{\bf B\times[\nabla\times B]}-\frac{{\bf\nabla}P_{e}}{en_{e}}.$
In presence of thermodiffusion and external electrical field $\bf E^{{}^{\prime\prime}}$ we obtain, after applying the curl procedure to the equation (20), and using (2), the following equation at $P_{e}=n_{e}kT$:
$\frac{\partial{\bf B}}{\partial t}={\bf\nabla}\times[{\bf v\times B}]+c{\bf\nabla\times E^{{}^{\prime\prime}}}+c\nabla\biggl(\frac{\lambda_{0}}{\sigma_{0}}\biggr)\times{\nabla T}-\frac{ck}{en_{e}}[{\bf\nabla}n_{e}\times{\bf\nabla}T]$ (21)
@@
3.1 The Biermann battery
Two last terms in the upper line of Eq.(21) may create a seed magnetic field in the non- magnetized media. The first term is related to the thermo-diffusion, and the second one is connected with a current produces by a pressure gradient (barodiffusion). The seed magnetic field is created when the surfaces of the constant concentration $n_{e}$ don’t coincide with surfaces of constant temperature,
$\nabla n_{e}\times\nabla T\neq 0.$
These two terms are representing the Biermann battery [7, 8].
The Eq.(21) is valid for variable values of the parameters $\lambda_{0}$ and $\sigma_{0}$. At constant value $\lambda_{0}/\sigma_{0}=3k/2e=const$, in the fully ionized non-degenerate Lorenz plasma, the first temperature term is zero. At in this conditions thermo-diffusion is not participating in creation of the seed magnetic field. When this term is variable,
$\lambda_{0}/\sigma_{0}=f(\rho,T)$
the seed magnetic field is created by a thermo-diffusion at non-zero value of
@@ 3.1 The Biermann battery
In neutron star crust the electrons are strongly degenerated, and the conditions of the seed field creation are changed.
Let us consider the case of strongly degenerate, non-relativistic electrons. In this case the coefficients of conductivity and thermo-diffusion, calculated in [20], have the form
$\sigma_{0}^{d}=\frac{e^{2}n_{e}}{m_{e}}\tau_{d},\qquad\lambda_{0}^{d}=\frac{4\pi^{3}e}{(9\pi)^{1/3}}\frac{n_{e}^{1/3}k^{2}T}{h^{2}}\tau_{d}.$ (22)
Here
$\tau_{d}=\frac{3}{32\pi^{2}}\frac{h^{3}n_{e}}{Z^{2}e^{4}m_{e}n_{N}\Lambda},\qquad h\,\,\,\,{\mbox{ is the Planck constant}}.$ (23)
The ratio of transport coefficients is written as
$\frac{\lambda_{0}^{d}}{\sigma_{0}^{d}}=\frac{k}{e}\,\frac{4\pi^{3}}{(9\pi)^{1/3}}\frac{kTm_{e}}{h^{2}n_{e}^{2/3}}$ (24)
what may give a non-zero value $\nabla n_{e}\times\nabla T\neq 0.$
For strongly degenerate plasma the baro-diffusion term in (21) should be taken, using (7), with the equation of state in the form (see [20])
$P_{e}=2kT\big(\frac{2kTm_{e}}{h^{2}}\big)^{3/2}G_{5/2}(x_{0}),\quad n_{e}=2\big(\frac{2kTm_{e}}{h^{2}}\big)^{3/2}G_{3/2}(x_{0}).$ (25)
We obtain from [21],[16]
$x_{0}=\frac{\varepsilon_{fe}}{kT}=\frac{(3\pi^{2}n_{e})^{2/3}h^{2}}{8\pi^{2}m_{e}kT}\gg 1,\qquad\qquad$ (26)
$G_{5/2}(x_{0})\approx\frac{x_{0}^{5/2}}{\Gamma(7/2)}+\frac{\pi^{2}}{4}\frac{x_{0}^{1/2}}{\Gamma(5/2)},\quad G_{3/2}(x_{0})\approx\frac{x_{0}^{3/2}}{\Gamma(5/2)}+\frac{\pi^{2}}{12}\,\frac{x_{0}^{-1/2}}{\Gamma(3/2)},$
$P_{e}=\frac{(3\pi^{2})^{2/3}}{20\pi^{2}}\frac{h^{2}n_{e}^{5/3}}{m_{e}}\big(1+\frac{5\pi^{2}}{8x_{0}^{2}}\big)=P_{ed}\big(1+\frac{5\pi^{2}}{8x_{0}^{2}}\big).$ (27)
The creation of the seed magnetic field in the strongly degenerate plasma is described by the equation
$\frac{\partial{\bf B}}{\partial t}=c\nabla\biggl(\frac{\lambda_{0}^{d}}{\sigma_{0}^{d}}\biggr)\times{\nabla T}+\frac{c}{e}\,\nabla\times\biggl(\frac{{\bf\nabla}P_{e}}{n_{e}}\biggr)$ (28)
We obtain from (24) and (27)
$\nabla\biggl(\frac{\lambda_{0}^{d}}{\sigma_{0}^{d}}\biggr)=\frac{k}{e}\,\frac{4\pi^{3}}{(9\pi)^{1/3}}\frac{km_{e}}{h^{2}}\biggl(\frac{\nabla T}{n_{e}^{2/3}}-\frac{2}{3}\frac{T\nabla n_{e}}{n_{e}^{5/3}}\biggr),$ (30)
$\nabla\times\biggl(\frac{{\bf\nabla}P_{e}}{n_{e}}\biggr)\,=-\frac{\nabla n_{e}}{n_{e}^{2}}\times{\bf\nabla P_{e}}=-2\frac{P_{ed}}{n_{e}^{2}}\frac{5\pi^{2}}{8x_{0}^{2}T}\nabla n_{e}\times{\nabla T}.$ (31)
With account of (30),(31) we have from (28)
$\frac{\partial{\bf B}}{\partial t}=-\biggl[\frac{4\pi^{8/3}}{3^{2/3}}\frac{c}{e}\frac{k^{2}Tm_{e}}{h^{2}\,n_{e}^{5/3}}+\frac{8\,\pi^{7/3}}{3^{5/3}}\,\frac{c}{e}\,\frac{k^{2}Tm_{e}}{h^{2}n_{e}^{5/3}}\biggr](\nabla n_{e}\times{\nabla T})$ (32)
Using (27), this equation is rewritten as (see also [26])
$\frac{\partial{\bf B}}{\partial t}=-\biggl[40.7109+18.5312\biggr]\frac{c}{e}\frac{k^{2}Tm_{e}}{h^{2}\,n_{e}^{5/3}}(\nabla n_{e}\times{\nabla T})$ (33)
In (33) the left term in the brackets is related to baro-diffision, and the right one is for thermo-diffusion. The input of both terms here is comparable, with a constant ratio.
3.2 The system of MHD equations, related to magnetic field behavior